Is y=3x-8 parallel, perpendicular, or neither to 3x-y=-1
Answer by algebrahouse.com(1659)

"Is y = 3x - 8 parallel, perpendicular, or neither to 3x - y = -1"

Parallel lines have equal slopes.
Perpendicular lines have slopes which are negative reciprocals, Ex: 2/3 and -3/2

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

In y = 3x - 8
the slope is 3 {the number in front of the x}

In 3x - y = -1
-y = -3x - 1 {subtracted 3x from both sides to get into slope-intercept form}
y = 3x + 1 {divided both sides by -1}
the slope is 3 {the number in front of the x}

The slopes are the same in both equations, therefore
the lines are parallel.
www.algebrahouse.com